# A 13/6-Approximation for Strip Packing via the Bottom-Left Algorithm

### v1

@@ 3.4 The proof of Theorem 1
Consider the packing $\text{BL}(\mathcal{I}_{\mathcal{FQW}})$. If $\mathcal{Q}=\varnothing$ or $y_{r}+h_{r}\leq h_{\text{max}}$ for all $r\in\mathcal{Q}$, then Lemma 5 implies that the BL algorithm is a $2$-approximation. Hence assume that $\mathcal{Q}\neq\varnothing$ and $\max\{y_{r}+h_{r}\mid r\in\mathcal{Q}\}>h_{\text{max}}$. The space $H^{\prime}$ defined in Lemma 8 is contained in $H_{a+1}\cup H_{L}\cup\cdots\cup H_{a+b+1}$ by Lemma 10. Thus, Lemma 678 and 9 imply that the space $H_{a+1}\cup H_{L}\cup\cdots\cup H_{a+b+1}$ is at least half occupied by rectangles. Furthermore, Theorem 17 states that the space $H_{a+2}\cup\cdots\cup H_{L}$ is at least $5/12$ occupied. This space has height at most $h_{\text{max}}$ by Lemma 11, because $y_{r_{L}}-h_{\text{max}}\leq(h_{\max}+h_{r_{B}})-h_{\text{max}}\leq h_{\text{max}}$. Thus the BL algorithm covers at least an area of
$\frac{1}{2}(h_{\text{BL}}-h_{\max})W+\frac{5}{12}h_{\max}W=\frac{1}{2}h_{\text{BL}}W-\frac{1}{12}h_{\max}W.$
@@
$h_{\text{OPT}}\geq\frac{1}{2}h_{\text{BL}}-\frac{1}{12}h_{\max}.$
@@
$h_{\text{BL}}\leq 2h_{\text{OPT}}+\frac{1}{6}h_{\max}\leq\frac{13}{6}h_{\text{OPT}}.\qed$
@@ 3.5.2 Lower bounds
The checkerboard instance of [1] gives a lower bound of $2$ on the approximation ratio of the BL algorithm when squares are packed in order of decreasing size. By extending this instance with a row of squares of size $4$, we obtain the same lower bound for the $\mathcal{FQW}$-ordering of squares or rectangles, as these size-$4$ squares constitute the set $\mathcal{F}$.
@@
Figure 4: The BL packing of an instance following the $\mathcal{FQW}$-ordering, where $\mathcal{F}=\{(w,h+1),(w+1,h)\}$,  $\mathcal{Q}=\{(w+1,h),(w+1,1)(w+1,h)\}$ and $\mathcal{W}=\varnothing$.
Consider the BL packing in $\mathcal{FQW}$-order for an instance of strip width $3w$ consisting of one rectangle of size $(w,h+1)$, three rectangles of size $(w+1,h)$, and one of size $(w+1,1)$ (Figure 4). As $h,w\to\infty$, the combined space $H^{\prime}=H_{3}\cup H_{6}$ is half occupied, while $H_{4}$ is only one-third occupied. The total height of $H_{3}\cup H_{6}$ approaches $h_{\text{OPT}}$, and the height of $H_{4}$ approaches $\frac{1}{2}h_{\text{OPT}}$. Using the area bound this implies that
@@
This example shows that narrowing the gap in the approximation ratio of the Bottom-Left Algorithm with $\mathcal{FQW}$-ordering requires new techniques that analyze the region $H_{a+2}\cup\cdots\cup H_{L}$ jointly with the other regions. In particular, we have
@@ 4 Conclusion
Determining the exact approximation ratio of the BL algorithm remains an open and intriguing problem. There are instances for which no ordering of rectangles can achieve a ratio better than $4/3-\varepsilon$ for the BL algorithm [11], even when all rectangles are squares. Closing the gap between $4/3-\varepsilon$ and $13/6$ is interesting, because of the possibility for the BL algorithm to surpass the current best-known $(5/3+\varepsilon)$-approximation for Strip Packing [9]. Furthermore, for the special case of Square Strip Packing no approximation algorithm with ratio better than $5/3+\varepsilon$ is known, this also motivates narrowing the gap between $4/3-\varepsilon$ and $2$ for the BL algorithm when only dealing with squares.

### v2

@@ 3.4 The proof of Theorem 1
Consider the packing $\text{BL}(\mathcal{I}_{\mathcal{FQW}})$. If $\mathcal{Q}=\varnothing$ or $y_{r}+h_{r}\leq h_{\text{max}}$ for all $r\in\mathcal{Q}$, then Lemma 5 implies that the BL algorithm is a $2$-approximation. Hence assume that $\mathcal{Q}\neq\varnothing$ and $\max\{y_{r}+h_{r}\mid r\in\mathcal{Q}\}>h_{\text{max}}$. The space $H^{\prime}$ defined in Lemma 8 is contained in $H_{a+1}\cup H_{L}\cup\cdots\cup H_{a+b+1}$ by Lemma 10. Thus, Lemma 678 and 9 imply that the space $H_{a+1}\cup H_{L}\cup\cdots\cup H_{a+b+1}$ is at least half occupied by rectangles. Furthermore, for $h=y_{r_{L}}-y_{r_{a+1}}$, Theorem 17 states that the space $H_{a+2}\cup\cdots\cup H_{L}$ is at least $\frac{1}{2}hW-\frac{1}{12}h_{\text{OPT}}W$ occupied. Thus the BL algorithm covers at least an area of
$\frac{1}{2}(h_{\text{BL}}-h)W+\frac{1}{2}hW-\frac{1}{12}h_{\text{OPT}}W=\frac{1}{2}h_{\text{BL}}W-\frac{1}{12}h_{\text{OPT}}W.$
@@
$h_{\text{OPT}}\geq\frac{1}{2}h_{\text{BL}}-\frac{1}{12}h_{\text{OPT}}.$
@@
$h_{\text{BL}}\leq\frac{13}{6}h_{\text{OPT}}.$
@@ 3.5.2 Lower bounds
The checkerboard instance of [1] gives a lower bound of $2$ on the approximation ratio of the BL algorithm when squares are packed in order of decreasing size. The same lower bound holds for the $\mathcal{FQW}$-ordering as the squares of size $2-i\varepsilon$ constitute the set $\mathcal{F}$.
@@
Figure 4: The BL packing of an instance following the $\mathcal{FQW}$-ordering on a strip of width $W=3w$ with rectangles $\mathcal{F}=\{(w,h+1),(w+1,h)\}$,  $\mathcal{Q}=\{(w+1,h),(w+1,1)(w+1,h)\}$ and $\mathcal{W}=\varnothing$.
Consider the BL packing in $\mathcal{FQW}$-order for an instance with strip width $3w$ consisting of one rectangle of size $(w,h+1)$, three rectangles of size $(w+1,h)$, and one of size $(w+1,1)$ (see Figure 4). As $h,w\to\infty$, the combined space $H^{\prime}=H_{3}\cup H_{6}$ is half occupied, while $H_{4}$ is only one-third occupied. The total height of $H_{3}\cup H_{6}$ approaches $h_{\text{OPT}}$, and the height of $H_{4}$ approaches $\frac{1}{2}h_{\text{OPT}}$. Using the area bound this implies that
@@
This example shows that narrowing the gap in the approximation ratio of the Bottom-Left Algorithm with $\mathcal{FQW}$-ordering requires new techniques that analyze the union of the regions $H_{a+2}\cup\cdots\cup H_{L}$ jointly with the other regions. In particular, we have
@@ 4 Conclusion
Determining the exact approximation ratio of the BL algorithm remains an open and intriguing problem. There are instances for which no ordering of rectangles can achieve a ratio better than $4/3-\varepsilon$ for the BL algorithm [11], even when all rectangles are squares. Closing the gap between $4/3-\varepsilon$ and $13/6$ is interesting, because of the possibility for the BL algorithm to surpass the current best-known $(5/3+\varepsilon)$-approximation for Strip Packing [9]. Furthermore, for the special case of Square Strip Packing no approximation algorithm with ratio better than $5/3+\varepsilon$ is known. This also motivates narrowing the gap between $4/3-\varepsilon$ and $2$ for the BL algorithm when only dealing with squares.
5 Appendix
In this appendix we show that the Bottom-Left Algorithm has unbounded approximation ratio for a large class of rectangle orderings, including those based on area or diagonal length. In particular, we construct a parameterized lower bound instance and prove that, for any ordering determined by a function $f$ satisfying certain conditions, the BL algorithm achieves a corresponding lower bound on its approximation factor. After establishing this lower bound theorem, we illustrate its effectiveness in showing that many orderings have unbounded approximation ratio. For example, this occurs when rectangles are placed according to their area, diagonal length, or any linear combination $aw_{r}+bh_{r}$ of their width and height, where $a,b>0$.
Theorem 18.
Let $f\colon(0,W]\times\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ be a function. Suppose that there exists a sequence $h_{1}<\cdots<h_{k}$ in $\mathbb{R}_{>0}$ and $y\in\mathbb{R}_{>0}$ such that the inverses $(f_{h_{i}})^{-1}(y)$ exist. And assume that for $W=\sum_{i=1}^{k}(f_{h_{i}})^{-1}(y)$, the inverse $(f_{W})^{-1}(y)$ also exists. Then the approximation ratio of the Bottom-Left Algorithm when placing rectangles in any order of $f$-value is at least
$\frac{h_{\text{BL}}^{f}}{h_{\text{OPT}}}\geq\frac{\sum_{i=1}^{k}h_{i}+k\cdot(f_{W})^{-1}(y)}{h_{k}+k\cdot(f_{W})^{-1}(y)}.$
Proof.
Consider a strip of width $W=\sum_{i=1}^{k}(f_{h_{i}})^{-1}(y)$ together with the rectangles $r_{i}=((f_{h_{i}})^{-1}(y),h_{i})$ and $q_{i}=(W,(f_{W})^{-1}(y))$ for $1\leq i\leq k$. The $f$-value of any of these rectangles equals $y$, hence for any ordering of the rectangles based solely on the $f$-values, the BL algorithm cannot distinguish these rectangles. Therefore, consider the ordering $r_{1},q_{1},\dots,r_{k},q_{k}$ that alternatingly places $r$ and $q$ rectangles by increasing $i$ index. As the height of the rectangles $r_{1},\dots,r_{k}$ is strictly increasing, and the rectangles $q_{i}$ have width $W$, it follows that $r_{i}$ is placed on top of $q_{i-1}$, and $q_{i}$ is placed on top of $r_{i}$. As all the rectangles are placed on top of each other, it follows that the height of the BL packing is the sum of the heights of all the rectangles.
An optimum packing places the rectangles $q_{1},\dots,q_{k}$ on top of each other, followed by placing the rectangles $r_{1},\dots,r_{k}$ next to each other, which is possible as the sum of their widths is exactly the width of the strip $W$. Thus the height of an optimum packing equals the sum of the height of the rectangles $q_{1},\dots,q_{k}$ and the height $h_{k}$ as this is the largest of the height amongst the rectangles $r_{1},\dots,r_{k}$. ∎
The first consequence of Theorem 18 is that ordering rectangles by area results in an unbounded approximation ratio of the BL algorithm.
Corollary 19.
The Bottom-Left Algorithm has unbounded approximation ratio when solely basing the order of the rectangles on the values of the function $f(w,h)=w^{p}h^{q}$ for $p,q>0$. In particular, this holds for ordering the rectangles by area.
Proof.
Define $\alpha=\frac{p}{q}$ and let $h_{i}=i^{\alpha}$ and $y=1$. Then $(f_{h_{i}})^{-1}(1)=\frac{1}{i}$, the strip width is $W=H_{k}$ (the $k$-th harmonic number) and $(f_{W})^{-1}(1)=\frac{1}{H_{k}^{\alpha}}$. By Theorem 18 we have the lower bound
$\frac{h_{\text{BL}}^{f}}{h_{\text{OPT}}}$ $\geq\frac{\sum_{i=1}^{k}i^{\alpha}+\frac{k}{H_{k}^{\alpha}}}{k^{\alpha}+\frac{k}{H_{k}^{\alpha}}}\geq\frac{H_{k}^{\alpha}\int_{0}^{k}x^{\alpha}\ dx+k}{H_{k}^{\alpha}\cdot k^{\alpha}+k}=\frac{\frac{1}{\alpha+1}H_{k}^{\alpha}\cdot k^{\alpha+1}+k}{H_{k}^{\alpha}\cdot k^{\alpha}+k}.$
This is unbounded for $k\to\infty$. For $p=q=1$ the function $f$ describes the area of a rectangle. ∎
The BL algorithm is a $3$-approximation when rectangles are ordered by decreasing width [1]. The following corollary is surprising, because it states that when we order the rectangles by decreasing width plus a tiny fraction of the height (e.g., by decreasing $w+10^{-100}\cdot h$ value), then the approximation ratio of the BL algorithm becomes unbounded.
Corollary 20.
The Bottom-Left Algorithm has unbounded approximation ratio when solely basing the order of the rectangles on the values of the function $f(w,h)=aw+bh$ with $a,b>0$.
Proof.
For $f(w,h)=aw+bh$, let $h_{i}=a(1+i\varepsilon)$$y=ab(1+k\varepsilon)$ and choose $\varepsilon=\frac{2}{k(k-1)}$. Then we have $(f_{h_{i}})^{-1}(y)=b(k-i)\varepsilon$, the width of the strip is $W=b$ and $(f_{W})^{-1}(y)=ak\varepsilon$. By Theorem 18 a lower bound on the approximation ratio is
$\frac{h_{\text{BL}}^{f}}{h_{\text{OPT}}}$ $\geq\frac{\sum_{i=1}^{k}a(1+i\varepsilon)+k\cdot\frac{2a}{k-1}}{a(1+k\varepsilon)+k\cdot\frac{2a}{k-1}}=\frac{k+\varepsilon\frac{k(k+1)}{2}+\frac{2k}{k-1}}{1+\frac{2}{k-1}+\frac{2k}{k-1}}$
$=\frac{k(k-1)+k+1+2k}{k-1+2+2k}=\frac{k^{2}+2k+1}{3k+1}.$
This is unbounded for $k\to\infty$. ∎
Corollary 21.
The Bottom-Left Algorithm has unbounded approximation ratio when solely basing the order of the rectangles on the length of their diagonal.
Proof.
For $f(w,h)=\sqrt{w^{2}+h^{2}}$, define $h_{i}=\sqrt{1+i\varepsilon}$$y=\sqrt{1+k\varepsilon}$ and choose $\varepsilon=(\sum_{i=1}^{k-1}\sqrt{i})^{-2}$. Then we have $(f_{h_{i}})^{-1}(y)=\sqrt{(k-i)\varepsilon}$, the width of the strip is $W=\sqrt{\varepsilon}\sum_{i=1}^{k}\sqrt{k-i}=1$ and $(f_{W})^{-1}(y)=\sqrt{k\varepsilon}$. By Theorem 18 it follows that
$\frac{h_{\text{BL}}^{f}}{h_{\text{OPT}}}$ $\geq\frac{\sum_{i=1}^{k}\sqrt{1+i\varepsilon}+k\cdot\sqrt{k\varepsilon}}{\sqrt{1+k\varepsilon}+k\cdot\sqrt{k\varepsilon}}\geq\frac{k\sqrt{1}+0}{\sqrt{1+1/k}+1/\sqrt{k}}=\frac{k\sqrt{k}}{\sqrt{k+1}+1}$
because $\varepsilon\leq k^{-2}$. This is unbounded for $k\to\infty$. ∎